### v1

# On Weighted Arboricity: Conductance–Resistance Bounds and Monoid Structure

Abstract
We study a conductance-weighted Nash–Williams density for a finite simple undirected graph $G=(V,E,c)$ with a conductance assignment $c:E\to[0,\infty)$:
$\mathcal{A}_{c}(G):=\max\bigl\{D_{c}(H):H\subseteq G\text{ connected},|V(H)|\geq 2\bigr\},\qquad D_{c}(H):=\frac{\sum_{e\in E(H)}c(e)}{|V(H)|-1}.$
This functional reduces to the classical Nash–Williams density when $c\equiv 1$, is isomorphism invariant, monotone under subgraphs and edge additions, positively homogeneous, and convex. We prove sharp global bounds
$\max_{e\in E}c(e)\leq\mathcal{A}_{c}(G)\leq\sum_{e\in E}c(e),$
with attainment by some connected subgraph. On the analytic side, we introduce a local variant and derive conductance–resistance inequalities using effective resistances in the ambient network. If $R_{G}(e)$ denotes the effective resistance between the endpoints of $e$ in $G$, we show that every connected $H\subseteq G$ satisfies
$\sum_{e\in E(H)}c(e)\,R_{G}(e)\leq|V(H)|-1,$
which in turn yields the Cauchy–Schwarz inequality
$D_{c}(H)\leq\sqrt{\frac{\sum_{e\in E(H)}c(e)/R_{G}(e)}{|V(H)|-1}}$
and hence an explicit resistance-based upper bound on $\mathcal{A}_{c}(G)$. On the structural side, we describe the algebraic behavior of $\mathcal{A}_{c}(G)$. We show that under edge-disjoint union, $\mathcal{A}_{c}(G)$ behaves as a max invariant: for a finite disjoint union of weighted graphs one has $\mathcal{A}_{c}(G)=\max_{i}\mathcal{A}_{c_{i}}(G_{i})$. In particular, disjoint union induces a commutative idempotent monoid structure at the level of isomorphism classes, with $\mathcal{A}_{c}(G)$ idempotent with respect to this operation.
Keywords. Weighted arboricity, Nash–Williams density, conductance, effective resistance, electrical networks, spanning trees, monoids.
MSC (2020). Primary 05C05, 05C42; Secondary 05C21, 05C22, 90C35.
1 Introduction
The arboricity of a graph $G$, defined as the minimum number of forests whose union covers every edge of $G$, admits a beautiful min-max characterisation due to Nash–Williams [1]: the arboricity is the ceiling of the maximum, over all connected subgraphs $H$ with $|V(H)|\geq 2$, of |E(H)|/(|V(H)|1)|E(H)|/(|V(H)|-1). This density interpretation has proven to be useful in structural and algorithmic graph theory [2, 3, 5].
Weighted analogues of arboricity have been studied in several directions. When edges carry positive integer weights, Gabow and Westermann [2] gave min-max theorems for the minimum number of forests needed to cover each edge at least its weight. Real-weighted versions, especially with weights interpreted as capacities or costs, appear in multicommodity flow and congestion minimization [4]. However, the natural conductance-weighted analogue, where the density numerator is the sum of conductances (reciprocal resistances) rather than cardinality or integer weights, has received surprisingly little direct attention.
Independently, the language of electrical networks has become a standard tool for studying graph densities. Foster’s theorem [6] and Rayleigh’s monotonicity principle imply that the effective resistance $R_{G}(u,v)$ between any two vertices is at most the reciprocal of the number of edge-disjoint paths, and this perspective has been used to bound spanning-tree densities and Laplacian matrix eigenvalues [8, 7]. Despite the close formal analogy between conductance-weighted spanning-tree counts and effective resistance, no systematic study of the resulting functional appears in the literature.
The purpose of this article is to fill this gap. For a finite undirected graph $G$ equipped with nonnegative edge conductances $c:E\to[0,\infty)$, we define the conductance-weighted arboricity
$\mathcal{A}_{c}(G):=\max\bigl\{D_{c}(H):H\subseteq G\text{ connected},|V(H)|\geq 2\bigr\},\qquad D_{c}(H):=\frac{\sum_{e\in E(H)}c(e)}{|V(H)|-1}.$
  1. 1.
    The functional $\mathcal{A}_{c}(G)$ inherits all expected analytic properties from the classical case: isomorphism invariance, monotonicity under edge addition and weight increase, positive homogeneity, convexity, and sharp global bounds $\max c(e)\leq\mathcal{A}_{c}(G)\leq\sum c(e)$, with attainment (Section 2).
  2. 2.
    Every connected subgraph $H$ satisfies $\sum_{e\in E(H)}c(e)R_{G}(e)\leq|V(H)|-1$, yielding an explicit resistance-based upper bound on $\mathcal{A}_{c}(G)$ (Section 3).
  3. 3.
    Under edge-disjoint union, $\mathcal{A}_{c}(G)$ behaves like a max invariant: $\mathcal{A}_{c}(G\sqcup H)=\max\{\mathcal{A}_{c}(G),\mathcal{A}_{c}(H)\}$. This induces a commutative idempotent monoid structure on the set of conductance-weighted isomorphism classes (Section 4).
While the basic analytic properties in Section 2 follow the standard approach for graph density and offer no surprises, the resistance inequality in Section 3 and especially the max-monoid observation in Section 4 appear to be new. In particular, we are unaware of any prior reference that records the idempotent monoid structure for arboricity (weighted or unweighted) under disjoint union, despite its similarity to the well-known behavior of girth, independence number, and odd-girth.
The rest of the paper is organised as follows. Section 2 recalls definitions and proves the basic properties. Section 3 introduces a local variant and derives the conductance–resistance inequalities. Section 4 establishes the monoid structure and concludes with remarks on possible functorial extensions.
2 Weighted Arboricity: Definitions and Basic Properties
2.1 Setup
Throughout, $G=(V,E,c)$ denotes a finite, simple, undirected graph with conductance assignment $c:E\to[0,\infty)$. For a subgraph $H\subseteq G$, we denote its vertex and edge sets by $V(H)$ and $E(H)$. Unless explicitly stated otherwise, all subgraphs considered in maximizations are assumed to be connected and satisfy $|V(H)|\geq 2$.
Definition 2.1 (Weighted density and weighted arboricity).
For any connected subgraph $H\subseteq G$ with $|V(H)|\geq 2$, define the weighted density
$D_{c}(H)\;:=\;\frac{\sum_{e\in E(H)}c(e)}{|V(H)|-1}.$
The weighted arboricity of $(G,c)$ is
$\mathcal{A}_{c}(G)\;:=\;\max\Bigl\{D_{c}(H)\;:\;H\subseteq G,\ H\text{ connected},\ |V(H)|\geq 2\Bigr\}.$
Remark 2.2 (Reduction from disconnected subgraphs).
If one allows disconnected $H$ and replaces the denominator by $|V(H)|-\omega(H)$, where $\omega(H)$ is the number of connected components of $H$, then the resulting maximization equals the one over connected $H$. Indeed, if $H=\bigsqcup_{i}H_{i}$ with each $H_{i}$ connected and $|V(H_{i})|\geq 2$, then
$\frac{\sum_{i}\sum_{e\in E(H_{i})}c(e)}{\sum_{i}\bigl(|V(H_{i})|-1\bigr)}$
is a weighted average of $\{D_{c}(H_{i})\}$ with weights $|V(H_{i})|-1\geq 1$, and hence does not exceed $\max_{i}D_{c}(H_{i})$. Therefore it suffices to optimize over connected subgraphs.
2.2 Basic properties
We write $n:=|V(G)|$, $m:=|E(G)|$, $c_{\max}:=\max_{e\in E(G)}c(e)$, and $C(G):=\sum_{e\in E(G)}c(e)$.
Proposition 2.3 (Reduction to the classical density).
If $c\equiv 1$ on $E(G)$, then
$\mathcal{A}_{c}(G)\;=\;\max_{H\subseteq G\text{ conn}}\frac{|E(H)|}{|V(H)|-1},$
the Nash–Williams density. In particular, $\lceil\mathcal{A}_{c}(G)\rceil$ equals the (unweighted) arboricity.
Proof.
With $c\equiv 1$, we have $\sum_{e\in E(H)}c(e)=|E(H)|$, so $D_{c}(H)=|E(H)|/(|V(H)|-1)$. Taking maxima yields the claim. ∎
Proposition 2.4 (Isomorphism invariance).
If $\varphi:G\to G^{\prime}$ is a graph isomorphism and $c^{\prime}$ is transported by $c^{\prime}(\varphi(e)):=c(e)$, then $\mathcal{A}_{c}(G)=\mathcal{A}_{c^{\prime}}(G^{\prime})$.
Proof.
The map $\varphi$ bijects connected subgraphs, preserves vertex and edge counts, and thus preserves each $D_{c}(H)$. Taking maxima over corresponding subgraphs gives the result. ∎
Proposition 2.5 (Monotonicity).
If $G^{\prime}$ is obtained from $G$ by adding edges (with nonnegative weights), then $\mathcal{A}_{c}(G^{\prime})\geq\mathcal{A}_{c}(G)$. If $c^{\prime}\geq c$ pointwise on $E(G)$, then $\mathcal{A}_{c^{\prime}}(G)\geq\mathcal{A}_{c}(G)$.
Proof.
Adding edges enlarges the feasible family of $H$; $D_{c}$ on previously feasible $H$ is unchanged. For weight increases, for each fixed $H$, $\sum_{e\in E(H)}c^{\prime}(e)\geq\sum_{e\in E(H)}c(e)$, hence $D_{c^{\prime}}(H)\geq D_{c}(H)$. Taking maxima yields the claim. ∎
Proposition 2.6 (Subgraph monotonicity).
If $F\subseteq G$ is any subgraph (with conductances given by restriction of $c$), then $\mathcal{A}_{c}(F)\leq\mathcal{A}_{c}(G)$.
Proof.
Every connected $H\subseteq F$ is a connected subgraph of $G$ with the same $D_{c}(H)$. Hence $D_{c}(H)\leq\mathcal{A}_{c}(G)$ for all such $H$. Taking the maximum over $H\subseteq F$ gives $\mathcal{A}_{c}(F)\leq\mathcal{A}_{c}(G)$. ∎
Proposition 2.7 (Positive homogeneity and convexity).
For any $\alpha\geq 0$, $\mathcal{A}_{\alpha c}(G)=\alpha\,\mathcal{A}_{c}(G)$. Moreover, for any conductance assignments $c^{(1)},c^{(2)}$ and $\lambda\in[0,1]$,
$\mathcal{A}_{\lambda c^{(1)}+(1-\lambda)c^{(2)}}(G)\ \leq\ \lambda\,\mathcal{A}_{c^{(1)}}(G)+(1-\lambda)\,\mathcal{A}_{c^{(2)}}(G).$
Proof.
For fixed $H$, $D_{\alpha c}(H)=\alpha D_{c}(H)$ and $D_{\lambda c^{(1)}+(1-\lambda)c^{(2)}}(H)=\lambda D_{c^{(1)}}(H)+(1-\lambda)D_{c^{(2)}}(H)$. Taking maxima over $H$ yields the statements. ∎
Proposition 2.8 (Global bounds and attainment).
The following hold:
  1. (i)
    $\mathcal{A}_{c}(G)\leq C(G)$.
  2. (ii)
    $\mathcal{A}_{c}(G)\geq c_{\max}$.
  3. (iii)
    If $G$ is connected and $c(e)\geq\underline{c}>0$ for all $e\in E(G)$, then $\mathcal{A}_{c}(G)\geq\underline{c}$.
  4. (iv)
    The maximum in the definition of $\mathcal{A}_{c}(G)$ is attained by some connected subgraph $H^{\star}\subseteq G$.
Proof.
(i) For any feasible $H$, $|V(H)|-1\geq 1$ and $\sum_{e\in E(H)}c(e)\leq C(G)$, so $D_{c}(H)\leq C(G)$.
(ii) For any $e=\{u,v\}$, the subgraph on $\{u,v\}$ with edge $e$ gives $D_{c}(H)=c(e)$; taking $e$ with maximal weight gives the bound.
(iii) For any spanning tree $T$ of $G$, $\mathcal{A}_{c}(G)\geq\big(\sum_{e\in E(T)}c(e)\big)/(|V(G)|-1)\geq\underline{c}$.
(iv) There are finitely many connected subgraphs of $G$; thus $\{D_{c}(H)\}$ has a maximum. ∎
Remark 2.9 (Classical families, unit weights).
If $c\equiv 1$ then:
  1. (a)
    If $G$ is a tree, then $\mathcal{A}_{c}(G)=1$.
  2. (b)
    For a cycle $C_{n}$ ($n\geq 3$), $\mathcal{A}_{c}(C_{n})=\dfrac{n}{\,n-1\,}$.
  3. (c)
    For the complete graph $K_{n}$ ($n\geq 2$), $\mathcal{A}_{c}(K_{n})=\dfrac{n}{2}$.
Proof.
(a) In a tree $H$, |E(H)|=|V(H)|1|E(H)|=|V(H)|-1, hence $D_{c}(H)=1$.
(b) Connected $H\subseteq C_{n}$ are paths or the whole cycle. The maximum |E(H)|/(|V(H)|1)|E(H)|/(|V(H)|-1) is n/(n1)n/(n-1).
(c) For $K_{n}$, the maximum of $\binom{k}{2}/(k-1)=k/2$ over $2\leq k\leq n$ is attained at $k=n$. ∎
3 Local Variants and Foster–Rayleigh Constraints
3.1 Local weighted arboricity
Definition 3.1 (Local weighted arboricity).
For $v\in V(G)$, define
$\mathcal{A}^{\mathrm{loc}}_{c}(v)\ :=\ \max\Bigl\{D_{c}(H):H\subseteq G\ \text{connected},\ v\in V(H),\ |V(H)|\geq 2\Bigr\}.$
Proposition 3.2 (Local–global equivalence).
With $\mathcal{A}_{c}(G)$ as above,
$\mathcal{A}_{c}(G)\;=\;\max_{v\in V(G)}\mathcal{A}^{\mathrm{loc}}_{c}(v).$
Proof.
Every connected $H$ contributing to $\mathcal{A}_{c}(G)$ contains some $v\in V(H)$ and hence contributes to $\mathcal{A}^{\mathrm{loc}}_{c}(v)$. Conversely, $\mathcal{A}^{\mathrm{loc}}_{c}(v)\leq\mathcal{A}_{c}(G)$ holds by definition for each $v$. Taking maxima gives equality. ∎
3.2 A local Foster inequality
Let $R_{e}^{(X)}$ denote the effective resistance between the endpoints of edge $e$ computed in the connected network $X$ with conductances. We use:
  • (Foster’s First Theorem [6]) For $X=(H,c|_{E(H)})$, $\sum_{e\in E(H)}c(e)\,R_{e}^{(X)}=|V(H)|-1$.
  • (Rayleigh monotonicity) Deleting edges or reducing conductances does not decrease effective resistance between any two vertices.
Theorem 3.3 (Local Foster inequality).
Let $H\subseteq G$ be any connected subgraph with inherited conductances. Then
$\sum_{e\in E(H)}c(e)\,R_{e}^{(G)}\ \leq\ |V(H)|-1.$
Proof.
Let $X:=(H,c|_{E(H)})$. For each $e\in E(H)$, Rayleigh yields $R_{e}^{(G)}\leq R_{e}^{(X)}$. Multiplying by $c(e)$ and summing over $e\in E(H)$ gives
$\sum_{e\in E(H)}c(e)\,R_{e}^{(G)}\ \leq\ \sum_{e\in E(H)}c(e)\,R_{e}^{(X)}\ =\ |V(H)|-1,$
using Foster’s First Theorem on $X$. ∎
Proposition 3.4 (Cauchy–Schwarz inequality).
For any connected $H\subseteq G$ with $|V(H)|\geq 2$,
$D_{c}(H)\;\leq\;\sqrt{\,\frac{\sum_{e\in E(H)}\dfrac{c(e)}{R_{e}^{(G)}}}{\,|V(H)|-1\,}}\,.$
Consequently,
$\mathcal{A}_{c}(G)\ \leq\ \max_{\begin{subarray}{c}H\subseteq G\\ H\ \mathrm{connected}\end{subarray}}\sqrt{\,\frac{\sum_{e\in E(H)}\dfrac{c(e)}{R_{e}^{(G)}}}{\,|V(H)|-1\,}}\,.$
Proof.
Let $a_{e}:=\sqrt{c(e)R_{e}^{(G)}}$ and $b_{e}:=\sqrt{c(e)/R_{e}^{(G)}}$. Then
$\sum_{e\in E(H)}c(e)=\sum a_{e}b_{e}\leq\big(\sum a_{e}^{2}\big)^{1/2}\big(\sum b_{e}^{2}\big)^{1/2}$
$=\sqrt{\big(\sum c(e)R_{e}^{(G)}\big)\big(\sum c(e)/R_{e}^{(G)}\big)}.$
Apply Theorem 3.3 and divide by |V(H)|1|V(H)|-1. ∎
$v_{1}$$v_{2}$$v_{3}$$v_{4}$c=2c=2c=1c=1c=3c=3c=1c=1c=2c=2Subgraph HHFor $H=\{v_{1},v_{2},v_{3}\}$: $|V(H)|=3$, $|E(H)|=3$ $\sum_{e\in E(H)}c(e)=2+1+3=6$ $D_{c}(H)=\frac{6}{3-1}=3$ Foster–Rayleigh: $\sum_{e\in E(H)}c(e)R_{G}(e)\leq|V(H)|-1$ Implies: $D_{c}(H)\leq\sqrt{\frac{\sum_{e\in E(H)}\frac{c(e)}{R_{G}(e)}}{|V(H)|-1}}$ Network $G$ with conductances $c:E\to[0,\infty)$
Figure 1: A small weighted network illustrating the Foster–Rayleigh inequality from Theorem $3.$ The highlighted subgraph $H$ contains vertices $\{v_{1},v_{2},v_{3}\}$ with three edges carrying conductances $c=2,1,3$ respectively. The inequality $\sum_{e\in E(H)}c(e)R_{G}(e)\leq|V(H)|-1$ provides an upper bound on the sum of conductance-weighted effective resistances, which in turn yields the Cauchy–Schwarz certificate for the weighted density $D_{c}(H)$.
Proposition 3.5 (Immediate local lower bounds).
For any connected $H\subseteq G$ with $|V(H)|\geq 2$,
$\mathcal{A}_{c}(G)\ \geq\ D_{c}(H)\ =\ \frac{\sum_{e\in E(H)}c(e)}{|V(H)|-1}\ \geq\ \max_{e\in E(H)}c(e).$
Proof.
The first inequality holds since $\mathcal{A}_{c}(G)$ is the maximum over connected $H$. The second follows from the feasible two-vertex subgraph containing a single edge $e$. ∎
3.3 Separation from totals or degrees
Remark 3.6 (Separation from totals/degree).
With unit weights $c\equiv 1$, the triangle $K_{3}$ and the $3$-edge star $K_{1,3}$ both satisfy $C(G)=3$ and $\Delta_{1}(G)=2$, yet
$\mathcal{A}_{1}(K_{3})=\frac{3}{2}\qquad\text{and}\qquad\mathcal{A}_{1}(K_{1,3})=1.$
Proof.
For $K_{3}$, the connected subgraph $H=K_{3}$ yields $D_{1}(H)=|E|/(|V|-1)=3/2$. For $K_{1,3}$, every connected subgraph is a tree, hence $D_{1}(H)=1$ for all such $H$, and so $\mathcal{A}_{1}(K_{1,3})=1$. ∎
4 Algebraic structure
In addition to its analytic and extremal properties, the weighted arboricity functional $\mathcal{A}_{c}(G)$ interacts in a clean way with natural constructions on weighted graphs. In this section we isolate one such feature: $\mathcal{A}_{c}(G)$ behaves predictably under edge-disjoint unions. At the level of isomorphism classes of weighted graphs, disjoint union induces a commutative monoid structure, and $\mathcal{A}_{c}(G)$ is idempotent with respect to this operation (Section 4.1). These observations justify viewing $\mathcal{A}_{c}(G)$ like a max invariant in structure with a simple algebraic profile.
4.1 Disjoint unions and idempotent max-structure
The functorial behavior of $\mathcal{A}_{c}(G)$ under weighted embeddings reflects its monotonicity with respect to enlarging graphs and increasing edge weights. A complementary structural feature appears when we combine graphs via edge-disjoint union. In that setting, $\mathcal{A}_{c}(G)$ is governed simply by the maximum of the component arboricities, and this induces a commutative idempotent monoid structure at the level of isomorphism classes. We record this next.
Definition 4.1 (Edge-disjoint union of weighted graphs).
Let $(G_{1},c_{1})$ and $(G_{2},c_{2})$ be finite weighted graphs. Assume that their vertex sets are disjoint and that there are no edges between $V(G_{1})$ and $V(G_{2})$; if necessary, we may relabel vertices to ensure this.
The edge-disjoint union of $(G_{1},c_{1})$ and $(G_{2},c_{2})$ is the weighted graph $(G,c)$ defined by
$G:=G_{1}\sqcup G_{2},\qquad V(G):=V(G_{1})\cup V(G_{2}),\quad E(G):=E(G_{1})\cup E(G_{2}),$
and
$c(e):=\begin{cases}c_{1}(e),&e\in E(G_{1}),\\[2.0pt] c_{2}(e),&e\in E(G_{2}).\end{cases}$
By construction, $G$ has exactly two connected components $G_{1}$ and $G_{2}$.
Proposition 4.2 (Weighted arboricity of an edge-disjoint union).
Let $(G_{1},c_{1})$ and $(G_{2},c_{2})$ be finite weighted graphs, and let $(G,c):=(G_{1},c_{1})\sqcup(G_{2},c_{2})$ be their edge-disjoint union as above. Then
$\mathcal{A}_{c}(G)\;=\;\max\bigl\{\mathcal{A}_{c_{1}}(G_{1}),\,\mathcal{A}_{c_{2}}(G_{2})\bigr\}.$
More generally, if $(G,c)$ is the edge-disjoint union of finitely many weighted graphs $(G_{i},c_{i})$, $1\leq i\leq k$, then
$\mathcal{A}_{c}(G)\;=\;\max_{1\leq i\leq k}\mathcal{A}_{c_{i}}(G_{i}).$
[…]
Figure 2: Disjoint union as a max-monoid: on graphs the operation is $\sqcup$, on values it is $\max$, and $\mathcal{A}_{c}$ is a homomorphism $(\mathcal{G},\sqcup)\to(\mathbb{R}_{\geq 0},\max)$.
Proof.
We first prove the two-component case.
Let $(G,c)=(G_{1},c_{1})\sqcup(G_{2},c_{2})$. By definition,
$\mathcal{A}_{c}(G)=\max\Bigl\{D_{c}(H)\,:\,H\subseteq G\ \text{connected},\ |V(H)|\geq 2\Bigr\}.$
Step 1: Any connected subgraph lies in a single component. Let $H\subseteq G$ be a connected subgraph with $|V(H)|\geq 2$. Since there are no edges between $V(G_{1})$ and $V(G_{2})$, any path in $G$ lies entirely inside $G_{1}$ or entirely inside $G_{2}$. If $V(H)$ met both $V(G_{1})$ and $V(G_{2})$, then there would be no path inside $H$ joining a vertex from $V(G_{1})\cap V(H)$ to a vertex from $V(G_{2})\cap V(H)$, contradicting the connectedness of $H$. Therefore
$H\subseteq G_{1}\quad\text{or}\quad H\subseteq G_{2}.$
Step 2: Densities agree with the component densities. Suppose $H\subseteq G_{i}$ for some $i\in\{1,2\}$. Then, by the definition of $c$,
$\sum_{e\in E(H)}c(e)=\sum_{e\in E(H)}c_{i}(e).$
Since $V(H)$ is the same regardless of whether we regard $H$ as a subgraph of $G$ or of $G_{i}$, it follows that
$D_{c}(H)=\frac{\sum_{e\in E(H)}c(e)}{|V(H)|-1}=\frac{\sum_{e\in E(H)}c_{i}(e)}{|V(H)|-1}=D_{c_{i}}(H).$
Step 3: Maximization splits over components. Using Step 1 and Step 2, we can rewrite the definition of $\mathcal{A}_{c}(G)$ as
$\mathcal{A}_{c}(G)$ $=\max\Bigl\{D_{c}(H)\,:\,H\subseteq G\ \text{connected},\ |V(H)|\geq 2\Bigr\}$
$=\small\max\Bigl(\max\{D_{c}(H):H\subseteq G_{1}\ \text{connected},\ |V(H)|\geq 2\},\,\max\{D_{c}(H):H\subseteq G_{2}\ \text{connected},\ |V(H)|\geq 2\}\Bigr)$
$=\max\bigl\{\mathcal{A}_{c_{1}}(G_{1}),\,\mathcal{A}_{c_{2}}(G_{2})\bigr\},$
which is the desired identity.
For a finite disjoint union of $k\geq 2$ components, we argue by induction on $k$. The case $k=2$ is exactly the argument above. Assume the statement holds for k1k-1 components. Write
$(G,c)=\Bigl(\bigsqcup_{i=1}^{k-1}(G_{i},c_{i})\Bigr)\sqcup(G_{k},c_{k})=:(H,d)\sqcup(G_{k},c_{k}).$
By the two-component case,
$\mathcal{A}_{c}(G)=\max\bigl\{\mathcal{A}_{d}(H),\,\mathcal{A}_{c_{k}}(G_{k})\bigr\}.$
By the induction hypothesis applied to $(H,d)$,
$\mathcal{A}_{d}(H)=\max_{1\leq i\leq k-1}\mathcal{A}_{c_{i}}(G_{i}).$
Combining these gives
$\mathcal{A}_{c}(G)=\max\Bigl(\max_{1\leq i\leq k-1}\mathcal{A}_{c_{i}}(G_{i}),\,\mathcal{A}_{c_{k}}(G_{k})\Bigr)=\max_{1\leq i\leq k}\mathcal{A}_{c_{i}}(G_{i}),$
completing the induction and the proof. ∎
Corollary 4.3 (Idempotent behavior under disjoint union).
Let $\mathcal{G}$ denote the set of isomorphism classes of finite weighted graphs, and let $\sqcup$ denote edge-disjoint union on representatives. Then:
  1. 1.
    $(\mathcal{G},\sqcup)$ is a commutative monoid with identity element given by the edgeless graph (with arbitrary vertex set and zero weights).
  2. 2.
    For any $G,c]\in\mathcal{G}$,
    $\mathcal{A}_{c}(G\sqcup G)=\max\{\mathcal{A}_{c}(G),\mathcal{A}_{c}(G)\}=\mathcal{A}_{c}(G),$
    so $\mathcal{A}_{c}(G)$ is idempotent with respect to disjoint union at the level of its values.
Acknowledgments
The author thanks Dr. Lon Mitchell, PhD, colleagues, and mentors at Eastern Michigan University for helpful comments on earlier drafts. Any errors are the author’s alone.

### v2

@@
 Abstract
-We study a conductance-weighted Nash–Williams density for a finite simple undirected graph $G=(V,E,c)$ with a conductance assignment $c:E\to[0,\infty)$:
+We study a conductance-weighted arboricity for a finite simple undirected graph $G=(V,E,c)$ with a conductance assignment $c:E\to[0,\infty)$:
 $\mathcal{A}_{c}(G):=\max\bigl\{D_{c}(H):H\subseteq G\text{ connected},|V(H)|\geq 2\bigr\},\qquad D_{c}(H):=\frac{\sum_{e\in E(H)}c(e)}{|V(H)|-1}.$
-This functional reduces to the classical Nash–Williams density when $c\equiv 1$, is isomorphism invariant, monotone under subgraphs and edge additions, positively homogeneous, and convex.
+This functional reduces to the classical Nash–Williams identity when $c\equiv 1$, is isomorphism invariant, monotone under subgraphs and edge additions, positively homogeneous, and convex.
 We prove sharp global bounds
@@
 On the analytic side, we introduce a local variant and derive conductance–resistance inequalities using effective resistances in the ambient network.
-If $R_{G}(e)$ denotes the effective resistance between the endpoints of $e$ in $G$, we show that every connected $H\subseteq G$ satisfies
-$\sum_{e\in E(H)}c(e)\,R_{G}(e)\leq|V(H)|-1,$
-which in turn yields the Cauchy–Schwarz inequality
-$D_{c}(H)\leq\sqrt{\frac{\sum_{e\in E(H)}c(e)/R_{G}(e)}{|V(H)|-1}}$
-and hence an explicit resistance-based upper bound on $\mathcal{A}_{c}(G)$.
+If $R^{G}_{\text{eff}(e)}$ denotes the effective resistance between the endpoints of $e$ in $G$, we show that every connected $H\subseteq G$ satisfies
+$\sum_{e\in E(H)}c(e)\,R^{(G)}_{\text{ eff }(e)}\leq|V(H)|-1,$
+which in turn yields the upper bound
+$D_{c}(H)\leq\sqrt{\frac{\sum_{e\in E(H)}c(e)/R^{(G)}_{\text{ eff }(e)}}{|V(H)|-1}}$
+and hence an explicit effective resistance-based upper bound on $\mathcal{A}_{c}(G)$.
 On the structural side, we describe the algebraic behavior of $\mathcal{A}_{c}(G)$.
@@
 In particular, disjoint union induces a commutative idempotent monoid structure at the level of isomorphism classes, with $\mathcal{A}_{c}(G)$ idempotent with respect to this operation.
-Keywords. Weighted arboricity, Nash–Williams density, conductance, effective resistance, electrical networks, spanning trees, monoids.
+We also provide a computational exhibit on the hypercube family $Q_{d}$, including random conductance sampling, illustrating numerical evaluation of the resulting resistance-based bound.
+Keywords. Weighted arboricity, conductance, effective resistance, electrical networks, spanning trees, monoids.
 MSC (2020). Primary 05C05, 05C42; Secondary 05C21, 05C22, 90C35.
 1 Introduction
-The arboricity of a graph $G$, defined as the minimum number of forests whose union covers every edge of $G$, admits a beautiful min-max characterisation due to Nash–Williams [1]: the arboricity is the ceiling of the maximum, over all connected subgraphs $H$ with $|V(H)|\geq 2$, of |E(H)|/(|V(H)|1)|E(H)|/(|V(H)|-1).
+The arboricity of a graph $G$, defined as the minimum number of forests whose union covers every edge of $G$, admits a classic optimization characterization due to Nash–Williams [1]: the arboricity is the ceiling of the maximum, over all connected subgraphs $H$ with $|V(H)|\geq 2$, of |E(H)|/(|V(H)|1)|E(H)|/(|V(H)|-1).
 This density interpretation has proven to be useful in structural and algorithmic graph theory [2, 3, 4].
@@
 Independently, the language of electrical networks has become a standard tool for studying graph densities.
-Foster’s theorem [6] and Rayleigh’s monotonicity principle imply that the effective resistance $R_{G}(u,v)$ between any two vertices is at most the reciprocal of the number of edge-disjoint paths, and this perspective has been used to bound spanning-tree densities and Laplacian matrix eigenvalues [8, 7].
+Foster’s theorem [6] and Rayleigh’s monotonicity principle imply that the effective resistance $R^{(G)}_{\text{ eff }(e)}$ between any two vertices is at most the reciprocal of the number of edge-disjoint paths, and this perspective has been used to bound spanning-tree densities and Laplacian matrix eigenvalues [7, 8].
 Despite the close formal analogy between conductance-weighted spanning-tree counts and effective resistance, no systematic study of the resulting functional appears in the literature.
@@
   2. 2.
-    Every connected subgraph $H$ satisfies $\sum_{e\in E(H)}c(e)R_{G}(e)\leq|V(H)|-1$, yielding an explicit resistance-based upper bound on $\mathcal{A}_{c}(G)$ (Section 3).
+    Every connected subgraph $H$ satisfies $\sum_{e\in E(H)}c(e)R^{(G)}_{\text{ eff }(e)}\leq|V(H)|-1$, yielding an explicit resistance-based upper bound on $\mathcal{A}_{c}(G)$ (Section 4).
   3. 3.
@@
 In particular, we are unaware of any prior reference that records the idempotent monoid structure for arboricity (weighted or unweighted) under disjoint union, despite its similarity to the well-known behavior of girth, independence number, and odd-girth.
-The rest of the paper is organised as follows.
-Section 2 recalls definitions and proves the basic properties.
+The rest of the paper is organized as follows: Section 3 recalls definitions and proves the basic properties.
 Section 4 introduces a local variant and derives the conductance–resistance inequalities.
-Section 4 establishes the monoid structure and concludes with remarks on possible functorial extensions.
+Section 5 establishes the monoid structure and Section 6 explores a small computational exhibit on families of hypercubes of the form $Q_{d}$.
+2 Preliminaries and Notation
+Throughout, graphs are finite, simple, and undirected.
+We write $G=(V,E)$ with $n:=|V|$ and $m:=|E|$.
+For $S\subseteq V$, $G[S]$ denotes the induced subgraph on $S$, and for $H\subseteq G$ we write $V(H)$ and $E(H)$ for its vertex and edge sets.
+A subgraph $H$ is connected if it has a single connected component.
+When working with weighted graphs we equip $G$ with an edge conductance function $c:E\to(0,\infty)$ and write $(G,c)$; all weights are assumed nonnegative and are interpreted as conductances.
+Definitions of the weighted density and the conductance-weighted arboricity functional $A_{c}(G)$ (and related notation such as $D_{c}(H)$) are given in the following section and will be used consistently thereafter.
 3 Weighted Arboricity: Definitions and Basic Properties
 3.1 Setup
-Throughout, $G=(V,E,c)$ denotes a finite, simple, undirected graph with conductance assignment $c:E\to[0,\infty)$.
-For a subgraph $H\subseteq G$, we denote its vertex and edge sets by $V(H)$ and $E(H)$.
-Unless explicitly stated otherwise, all subgraphs considered in maximizations are assumed to be connected and satisfy $|V(H)|\geq 2$.
 Definition 3.1 (Weighted density and weighted arboricity).
@@
 3.2 Basic properties
-We write $n:=|V(G)|$, $m:=|E(G)|$, $c_{\max}:=\max_{e\in E(G)}c(e)$, and $C(G):=\sum_{e\in E(G)}c(e)$.
 Proposition 3.3 (Reduction to the classical density).
@@
 Proposition 3.8 (Global bounds and attainment).
-The following hold:
+Let $c_{\max}:=\max_{e\in E(G)}c(e)$, and $C(G):=\sum_{e\in E(G)}c(e)$. The following hold:
   1. (i)
@@
     For the complete graph $K_{n}$ ($n\geq 2$), $\mathcal{A}_{c}(K_{n})=\dfrac{n}{2}$.
+[…]
+Figure 1: Two weighted trees with edge conductances labeled.
+Since every connected subgraph of a tree is a subtree, the density $D_{c}(H)$ is an average of edge conductances over $H$, hence maximized by a single edge of largest conductance; therefore $\mathcal{A}_{c}(T)=\max_{e\in E(T)}c(e)$.
+$c\equiv 1,\qquad|V(K_{7})|=7,\qquad|E(K_{7})|=\binom{7}{2}=21,\qquad D_{1}(K_{7})=\frac{|E|}{|V|-1}=\frac{21}{6}=\frac{7}{2},\qquad\mathcal{A}_{1}(K_{7})=\frac{7}{2}.$
+Figure 2: Complete graph $K_{7}$ (unit conductances)
 Proof.
@@
 (b) Connected $H\subseteq C_{n}$ are paths or the whole cycle. The maximum |E(H)|/(|V(H)|1)|E(H)|/(|V(H)|-1) is n/(n1)n/(n-1).
-(c) For $K_{n}$, the maximum of $\binom{k}{2}/(k-1)=k/2$ over $2\leq k\leq n$ is attained at $k=n$. ∎
-3 Local Variants and Foster–Rayleigh Constraints
+(c) For $K_{n}$, the maximum of $\binom{k}{2}/(k-1)=k/2$ over $2\leq k\leq n$ is attained at $k=n$.
+4 Local Variants and Local Foster Constraints
 4.1 Local weighted arboricity
@@
 ∎
+4.2 Effective resistance via energy dissipation
+Here we provide a brief but fully rigorous characterization of effective resistances relying heavily on standard results from [9].
+So let $\bm{{v}}$ be a potential on the vertices $v\in V(G)$, $\bm{{W}}$ a diagonal weights matrix, and $\bm{{U}}$ a signed edge-vertex adjacency matrix.
+Also define a flow on the edges $e\in E(G)$ denoted $\bm{i}$}, with the external flow at vertices denoted $\bm{i_{ext}}$.
+Then define the Laplacian matrix, denoted $\bm{{L}}$, as
+$\bm{L:=U^{\top}WU}.$
+It satisfies the following conditions:
+  • $\bm{W}=\bm{R}^{-1}$ for $\bm{R}$, a resistances matrix, [^1] It is important to note that we are not yet discussing the formal object effective resistance, rather the simple resistances defined as inverse of their corresponding conductances.
+  • $\bm{i=WU{v}}$,
+  • $\bm{i_{ext}=U^{\top}i}$,
+  • $\bm{i_{ext}=Lv}$.
+[…]
+With the pieces in place, we define effective resistance $R^{(G)}_{\text{eff }(a,b)}$ as the resistance between some $a,b\in V(G)$ given by the whole network.
+This is useful for a standard but powerful characterization of effective resistance:
+Lemma 4.3 (Effective resistance).
+[…]
+Proof.
+Using $\bm{i_{ext}=Lv}$ and $\bm{v=i_{ext}L^{\dagger}v}$, we have
+[…]
+Thus we have two invariant weighted metrics which, while admitting different interpretation of edge weights, embed information locally at each edge and can thus be compared piecewise analytically in a formal way; such is the focus of this section, so we can now define one of the main contributions of this paper.
 4.3 A local Foster inequality
-Let $R_{e}^{(X)}$ denote the effective resistance between the endpoints of edge $e$ computed in the connected network $X$ with conductances. We use:
-  • (Foster’s First Theorem [6]) For $X=(H,c|_{E(H)})$, $\sum_{e\in E(H)}c(e)\,R_{e}^{(X)}=|V(H)|-1$.
+Let $R_{\text{eff}(e)}^{(X)}$ denote the effective resistance between the endpoints of edge $e$ computed in the connected network $X$ with conductances. We use:
+  • (Foster’s First Theorem [6]) For $X=(H,c|_{E(H)})$, $\sum_{e\in E(H)}c(e)\,R_{\text{eff}(e)}^{(X)}=|V(H)|-1$.
   • (Rayleigh monotonicity) Deleting edges or reducing conductances does not decrease effective resistance between any two vertices.
@@
 Let $H\subseteq G$ be any connected subgraph with inherited conductances. Then
-$\sum_{e\in E(H)}c(e)\,R_{e}^{(G)}\ \leq\ |V(H)|-1.$
+$\sum_{e\in E(H)}c(e)\,R_{\text{eff}(e)}^{(G)}\ \leq\ |V(H)|-1.$
 Proof.
-Let $X:=(H,c|_{E(H)})$. For each $e\in E(H)$, Rayleigh yields $R_{e}^{(G)}\leq R_{e}^{(X)}$. Multiplying by $c(e)$ and summing over $e\in E(H)$ gives
-$\sum_{e\in E(H)}c(e)\,R_{e}^{(G)}\ \leq\ \sum_{e\in E(H)}c(e)\,R_{e}^{(X)}\ =\ |V(H)|-1,$
+Let $X:=(H,c|_{E(H)})$. For each $e\in E(H)$, Rayleigh monotonicity yields $R_{\text{eff}(e)}^{(G)}\leq R_{\text{eff}(e)}^{(X)}$. Multiplying by $c(e)$ and summing over $e\in E(H)$ gives
+$\sum_{e\in E(H)}c(e)\,R_{\text{eff}(e)}^{(G)}\ \leq\ \sum_{e\in E(H)}c(e)\,R_{\text{eff}(e)}^{(X)}\ =\ |V(H)|-1,$
 using Foster’s First Theorem on $X$. ∎
-Proposition 3.4 (Cauchy–Schwarz inequality).
+Proposition 4.5 (An explicit upper bound).
 For any connected $H\subseteq G$ with $|V(H)|\geq 2$,
-$D_{c}(H)\;\leq\;\sqrt{\,\frac{\sum_{e\in E(H)}\dfrac{c(e)}{R_{e}^{(G)}}}{\,|V(H)|-1\,}}\,.$
+$D_{c}(H)\;\leq\;\sqrt{\,\frac{\sum_{e\in E(H)}\dfrac{c(e)}{R_{\text{eff}(e)}^{(G)}}}{\,|V(H)|-1\,}}\,.$
 Consequently,
-$\mathcal{A}_{c}(G)\ \leq\ \max_{\begin{subarray}{c}H\subseteq G\\ H\ \mathrm{connected}\end{subarray}}\sqrt{\,\frac{\sum_{e\in E(H)}\dfrac{c(e)}{R_{e}^{(G)}}}{\,|V(H)|-1\,}}\,.$
+$\mathcal{A}_{c}(G)\ \leq\ \max_{\begin{subarray}{c}H\subseteq G\\ H\ \mathrm{connected}\end{subarray}}\sqrt{\,\frac{\sum_{e\in E(H)}\dfrac{c(e)}{R_{\text{eff}(e)}^{(G)}}}{\,|V(H)|-1\,}}\,.$
 Proof.
-Let $a_{e}:=\sqrt{c(e)R_{e}^{(G)}}$ and $b_{e}:=\sqrt{c(e)/R_{e}^{(G)}}$. Then
-$\sum_{e\in E(H)}c(e)=\sum a_{e}b_{e}\leq\big(\sum a_{e}^{2}\big)^{1/2}\big(\sum b_{e}^{2}\big)^{1/2}$
-$=\sqrt{\big(\sum c(e)R_{e}^{(G)}\big)\big(\sum c(e)/R_{e}^{(G)}\big)}.$
-Apply Theorem 3.3 and divide by |V(H)|1|V(H)|-1. ∎
-$v_{1}$$v_{2}$$v_{3}$$v_{4}$c=2c=2c=1c=1c=3c=3c=1c=1c=2c=2Subgraph HHFor $H=\{v_{1},v_{2},v_{3}\}$: $|V(H)|=3$, $|E(H)|=3$ $\sum_{e\in E(H)}c(e)=2+1+3=6$ $D_{c}(H)=\frac{6}{3-1}=3$ Foster–Rayleigh: $\sum_{e\in E(H)}c(e)R_{G}(e)\leq|V(H)|-1$ Implies: $D_{c}(H)\leq\sqrt{\frac{\sum_{e\in E(H)}\frac{c(e)}{R_{G}(e)}}{|V(H)|-1}}$ Network $G$ with conductances $c:E\to[0,\infty)$
-Figure 1: A small weighted network illustrating the Foster–Rayleigh inequality from Theorem $3.$ The highlighted subgraph $H$ contains vertices $\{v_{1},v_{2},v_{3}\}$ with three edges carrying conductances $c=2,1,3$ respectively.
-The inequality $\sum_{e\in E(H)}c(e)R_{G}(e)\leq|V(H)|-1$ provides an upper bound on the sum of conductance-weighted effective resistances, which in turn yields the Cauchy–Schwarz certificate for the weighted density $D_{c}(H)$.
+Let $a_{e}:=\sqrt{c(e)R_{\text{eff}(e)}^{(G)}}$ and $b_{e}:=\sqrt{c(e)/R_{\text{eff}(e)}^{(G)}}$. Then
+$\sum_{e\in E(H)}c(e)=\sum a_{e}b_{e}\leq\big(\sum a_{e}^{2}\big)^{1/2}\big(\sum b_{e}^{2}\big)^{1/2}\text{ (by Cauchy--Schwarz) }$
+$=\sqrt{\big(\sum c(e)R_{\text{eff}(e)}^{(G)}\big)\big(\sum c(e)/R_{\text{eff}(e)}^{(G)}\big)}.$
+By Theorem 4.4, $\sum_{e\in E(H)}c(e)\,R_{\text{eff}(e)}^{(G)}\leq|V(H)|-1$, hence
+$\Bigl(\sum_{e\in E(H)}c(e)\Bigr)^{2}\leq(|V(H)|-1)\Bigl(\sum_{e\in E(H)}\frac{c(e)}{R^{(G)}_{\text{eff}(e)}}\Bigr).$
+Dividing by |V(H)|1|V(H)|-1,
+$\frac{\sum_{e\in E(H)}c(e)}{|V(H)|-1}=D_{c}(H)\leq\sqrt{\frac{\sum_{e\in E(H)}\frac{c(e)}{R^{(G)}_{\text{eff}(e)}}}{|V(H)|-1}}.$
+Thus take maxima over subgraphs to yield the result. ∎
+$v_{1}$$v_{2}$$v_{3}$$v_{4}$c=2c=2c=1c=1c=3c=3c=1c=1c=2c=2Subgraph HHFor $H=\{v_{1},v_{2},v_{3}\}$: $|V(H)|=3$, $|E(H)|=3$ $\sum_{e\in E(H)}c(e)=2+1+3=6$ $D_{c}(H)=\frac{6}{3-1}=3$ Local Foster: $\sum_{e\in E(H)}c(e)R^{(G)}_{\text{ eff }(e)}\leq|V(H)|-1$ Implies: $D_{c}(H)\leq\sqrt{\frac{\sum_{e\in E(H)}\frac{c(e)}{R^{(G)}_{\text{ eff }(e)}}}{|V(H)|-1}}$
+Figure 3: A small weighted network illustrating the Local Foster inequality from Theorem $4.$ The highlighted subgraph $H$ contains vertices $\{v_{1},v_{2},v_{3}\}$ with three edges carrying conductances $c=2,1,3$ respectively.
+The inequality $\sum_{e\in E(H)}c(e)R^{(G)}_{\text{ eff }(e)}\leq|V(H)|-1$ provides an upper bound on the sum of conductance-weighted effective resistances, which in turn yields the explicit upper bound for the weighted density $D_{c}(H)$.
 Proposition 4.6 (Immediate local lower bounds).
@@
 $\mathcal{A}_{1}(K_{3})=\frac{3}{2}\qquad\text{and}\qquad\mathcal{A}_{1}(K_{1,3})=1.$
+$K_{3}$
+$K_{1,3}$
+Figure 4: Two graphs with the same totals ($C(G)=3$) and maximum degree ($\Delta_{1}(G)=2$), but different $\mathcal{A}_{1}$.
 Proof.
@@
 ∎
+This is, of course, a somewhat trivial implication of the reduction to classical arboricity for $\mathcal{A}_{1}(G)$. Further results on heterogeneous $c$ are discussed in Section 6.
 5 Algebraic structure
@@
 […]
-Figure 2: Disjoint union as a max-monoid: on graphs the operation is $\sqcup$, on values it is $\max$, and $\mathcal{A}_{c}$ is a homomorphism $(\mathcal{G},\sqcup)\to(\mathbb{R}_{\geq 0},\max)$.
+Figure 5: Disjoint union as a monoid: $\mathcal{A}_{c}$ is a homomorphism $(\mathcal{G},\sqcup)\to(\mathbb{R}_{\geq 0},\max)$.
 Proof.
@@
     so $\mathcal{A}_{c}(G)$ is idempotent with respect to disjoint union at the level of its values.
+6 Computation
+This section gives a small computational exhibit for the numerical procedure used to generate the overlay figures comparing the baseline density $A_{1}(G)$ against the upper bound from Section 4, $\mathrm{CSBound}(H{=}G)$, in both the unit–conductance and random $c$ cases.
+All plots in this section were produced by the script scripts/make_figs.py, which outputs (i) an overlay plot of $A_{c}(G)$ and $\mathrm{CSBound}(H{=}G)$ versus $v=|V|$ for some common graph families, (ii) a tightness–ratio plot $\mathrm{CSBound}(H{=}G)/A_{c}(G)$ when defined, and (iii) a CSV summary of values and runtimes.
+While the aforementioned exhibits may be very useful for future analysis, their inclusion in this paper will be limited to one graph family for brevity.
+Accordingly, we restrict attention here to the hypercube family $Q_{d}$, which is highly structured (edge–transitive) and thus admits a clean, representative testbed for the bound without proliferating additional figures.
+6.1 Quantities plotted
+For a graph $G=(V,E)$ with $|V|=v\geq 2$, the script plots the unit–weight functional
+$A_{1}(G)\;=\;\max\Bigl\{\frac{|E(H)|}{|V(H)|-1}:\ H\subseteq G\ \text{is connected}\Bigr\}.$
+For very small graphs ($v\leq 12$ by default) this is computed by brute force enumeration over all connected induced vertex subsets.
+For larger graphs, the script returns $A_{1}(G)$ only in families where the exact value is known a priori without further optimization: for trees (paths and stars), $A_{1}(G)=1$; for complete graphs, cycles, complete bipartite graphs, and hypercubes, $A_{1}(G)=D_{1}(G)=|E|/(|V|-1)$.
+The explicit upper bound evaluated at $H=G$ in the unit conductance case comes from Proposition 4.5:
+$\mathrm{CSBound}(H{=}G)\;:=\;D_{c}(H=G)\;\leq\;\sqrt{\,\frac{\sum_{e\in E(H)}\dfrac{c(e)}{R_{\text{eff}(e)}^{(G)}}}{\,|V(H)|-1\,}}\,.$
+where $R^{(G)}_{\mathrm{eff}}$ is the effective resistance between the endpoints of $e$ in the ambient graph $G$.
+In the overlay figures we plot the pair $\bigl(A_{1}(G),\mathrm{CSBound}(H{=}G)\bigr)$ against $v=|V|$, and in the tightness figures we plot the ratio $\mathrm{CSBound}(H{=}G)/A_{1}(G)$ (when $A_{1}(G)>0$ and both quantities are finite).
+6.2 Computation on the hypercube family
+$Q_{2}$$Q_{3}$$Q_{4}$
+Figure 6: Hypercube family $Q_{d}$ for $d=2,3,4$
+We illustrate the computation of the upper bound $\mathrm{CSBound}(H{=}G)$ on the hypercube family $Q_{d}$.
+Since $Q_{d}$ is edge-transitive, all edges lie in one automorphism orbit; hence the ambient effective resistance $R_{\mathrm{eff}}^{(Q_{d})}(e)$ is the same for every edge $e\in E(Q_{d})$ [11, 12].
+Fix any edge $e_{0}=\{u,v\}$ and write
+$R_{d}\;:=\;R_{\mathrm{eff}(e_{0})}^{(Q_{d})}\;=\;R_{\mathrm{eff}(e)}^{(Q_{d})}\qquad\text{for all }e\in E(Q_{d}).$
+Hence for an arbitrary conductance assignment $c:E(Q_{d})\to[0,\infty)$,
+$\sum_{e\in E(Q_{d})}\frac{c(e)}{R_{\mathrm{eff}(e)}^{(Q_{d})}}\;=\;\frac{1}{R_{d}}\sum_{e\in E(Q_{d})}c(e).$
+To compute $R_{\mathrm{eff}(e)}^{(Q_{d})}$, the script follows from [13] and uses a grounded Laplacian solve: relabel vertices as $\{0,1,\dots,|V|-1\}$ with ground node 0, then form the reduced (grounded) Laplacian by deleting the row and column of the Laplacian corresponding to node 0, and for $u,v\in V(G)$, the endpoints of $e$, solve $L_{\mathrm{gr}}x=b$ with $b$ equal to +1+1 at $u$ and 1-1 at $v$ (with the grounded coordinate omitted).
+The effective resistance is then the resulting voltage drop $R_{\mathrm{eff}(e)}^{(Q_{d})}=|x_{u}-x_{v}|$ [13].
+6.3 Random conductance sampling
+To test robustness under heterogeneous weights, we generate i.i.d.
+random conductances on the hypercube edges from a two-point law supported on $\{1,\lambda\}$.
+Fix parameters $(\lambda,p)$ and, independently for each edge $e\in E(Q_{d})$, sample
+$c(e)=\begin{cases}\lambda,&\text{with probability }p,\\ 1,&\text{with probability }1-p.\end{cases}$
+For each dimension $d\in\{d_{\min},\dots,d_{\max}\}$ we draw reps independent conductance assignments and compute, for each draw, the overlay quantities $D_{c}(Q_{d})$ and $\mathrm{CSBound}(H{=}Q_{d})$, as well as the tightness ratio $\mathrm{CSBound}(H{=}Q_{d})/D_{c}(Q_{d})$.
+We summarize trial-to-trial variation in the overlay plots via shaded bands, and we report per-$d$ median ratios in the tightness figures.
+As a representative run, the command
+ --dist two_point --lambda 10 --p 0.5 --d_min 4 --d_max 10 --reps 25
+produces median tightness ratios decreasing with $d$, with total runtime on the order of one minute for the full sweep.
+Unless stated otherwise, we use $p=0.5$ and repeat the same experiment for $\lambda\in\{5,10,50\}$.
+6.4 Plots
+Figure 7: Densities with heterogeneous $c$ and $\lambda\in[1,5]$
+Figure 8: Tightness with heterogeneous $c$ and $\lambda\in[1,5]$
+Remark 6.1.
+Pleasantly, our properties from Section 3 show up in the asymptotic behavior of the $D_{c}(G)$ functional (prior to maximization over connected subgraphs), such as monotonicity.
+Additionally, the upper bound is tight at large $v$ (although these are all low-dimensional hypercube families.)
+Figure 9: Densities with heterogeneous $c$ and $\lambda\in[1,10]$
+Figure 10: Tightness with heterogeneous $c$ and $\lambda\in[1,10]$
+Figure 11: Densities with heterogeneous $c$ and $\lambda\in[1,50]$
+Figure 12: Tightness with heterogeneous $c$ and $\lambda\in[1,50]$
+Remark 6.2.
+Indeed, the bound is tight at asymptotically moderately large $v$; further research is needed to fully explore the computational complexity of different graph families.
+In short, the hypercube family provides a clean stress-test where symmetry allows us to reduce the evaluation of the certificate to a single grounded Laplacian solve (up to relabeling), while still permitting heterogeneous conductance assignments through random sampling.
+Te overlay figures indicate that $\mathrm{CSBound}(H{=}Q_{d})$ tracks the density functional $D_{c}(Q_{d})$ closely, and the tightness ratios remain near $1$ even as $|V(Q_{d})|=2^{d}$ grows (in particular, the tightness ratio never dips below 1 (Proposition 4.5)).
+These experiments suggest that the certificate is not merely an artifact of unit weights, but continues to behave stably under significant weight heterogeneity, motivating the broader family-wise computations and complexity discussion in future projects.
+7 Conclusion
+This paper introduced and studied a conductance–weighted analogue of arboricity built from the density functional
+$D_{c}(H)=\frac{\sum_{e\in E(H)}c(e)}{|V(H)|-1},$
+together with the associated extremal quantity obtained by maximizing over connected subgraphs.
+The guiding theme is that embedding conductances into the Nash–Williams framework produces an invariant that is both structurally similar to classical arboricity and also naturally tethered to effective resistance and electrical network geometry.
+On the theoretical side, we developed a usable upper bound for the extremal density in terms of effective resistances in the ambient graph, which is explicit once effective resistances are available.
+The resulting inequality behaves well under scaling, interfaces cleanly with classical families, and is computable in practice via standard linear algebra.
+A brief computational exhibit complements the theory.
+Using Python, we generated overlay and tightness-ratio plots comparing the baseline density to the explicit certificate across graph sizes.
+We emphasized the hypercube family $Q_{d}$, where edge–transitivity implies that all edges share the same effective resistance, so $\mathrm{CSBound}(H{=}G)$ can be evaluated from a single grounded–Laplacian solve per dimension.
+Both in the unit-conductance base case and under random two-point conductances, the certificate tracks the observed densities closely, and the tightness ratio trends toward $1$ as $d$ grows, supporting the view that the bound is sharp on highly symmetric families.
+There are two natural directions for future work.
+First, we pursue an economics application in which conductances encode intensity or reliability of interactions in a network (such as global trade or social contagion), so that the weighted arboricity functional measures a density that incorporates both combinatorial connectivity and resistance geometry.
+Second, on the purely mathematical side, the objects introduced here invite a more systematic analysis and algebraic formalization: finer structural properties of the extremal functional or limiting behavior under graph operations.
+We expect these two strains, applied network modeling and deeper structural theory, to reinforce each other with computation serving as a useful tool in that pursuit.
 Acknowledgments